### v1

# Finite-Sample Learning Control for LQR Over Unknown Lossy Channels

Abstract
This paper investigates the Linear Quadratic Regulator (LQR) problem over an unknown Bernoulli packet drop channel. The unknown packet drop probability is estimated using finite samples, then the estimated probability is used to design a formally equivalent optimal controller. If the estimation error is too large, the estimated controller cannot mean-square stabilize the system. For $n$-dimensional systems, the upper bound on the estimation error is provided to guarantee the stability of the closed-loop system. And we present an analytical expression for the gap between the performance of the estimated controller and the optimal performance. Next, we derive the upper bound on sample complexity for the stabilizability of the estimated controller. The tailored results with less conservatism are delivered for scalar systems and $n$-dimensional systems with invertible input matrix. Furthermore, a sufficient condition that does not depend on unknown channel information is provided to determine whether the estimated controller stabilizes the system with a certain probability. Finally, Numerical examples are used to demonstrate our results.
{IEEEkeywords}
LQR, Bernoulli packet drop, networked control systems, stabilizability threshold, sample complexity, performance analysis.
1 Introduction
Networked control systems (NCSs) connect plants and controllers in closed-loop through wireless communication networks. They offer many advantages and have been applied in various fields such as industrial automation and autonomous driving [1, 2, 3, 4, 5]. Meanwhile, NCSs may also introduce new challenges due to the vulnerability and openness of wireless communication [6]. Packet drop, as a common problem in wireless channels, disrupts the connection loop between different devices [14, 7, 8, 9, 10, 11, 13, 12]. The seminal work [9] firstly considers Kalman filtering with packet loss, where the loss of system observations follows a Bernoulli distribution. Subsequently, control problems over lossy channels were studied in [10, 11]. The work [10] investigates the LQR problem over Bernoulli packet loss channels, where packets in sensor-controller and controller-actuator channels may be lost. The work [11] extends LQR with packet drop to the Linear Quadratic Gaussian (LQG) problem and proves that the separation principle still holds under the TCP protocol. According to [10, 11], we know that the optimal controller over lossy channels depends on the channel information, i.e. the packet drop probability. Additionally, the works [13, 12, 14] present suboptimal controllers for NCSs with packet loss compensation mechanisms, which also rely on channel information.
However, obtaining channel information may involve high costs or even be impossible [15, 16]. Many studies have begun to exploit NCSs over unknown channels [17, 18, 16, 19, 20, 21, 22]. The work [19] considers control problems over multiple channels with unknown Bernoulli packet drop and employs Thompson sampling to learn loss probabilities, without restrictions on sample size. Yet, in practice, all data are finite. Novel perspectives based on finite sample learning and non-asymptotic analysis provide a more detailed description of the learning difficulty [23, 24]. The works [20, 21] consider a linear system with unknown Bernoulli packet drop, which is stable if the drop probability is sufficiently low and unstable if it is not. They use finite samples to estimate the unknown probability and then answer the system is stable or not on a confidence level. The key of these works is the Hoeffding’s inequality which establishes a bridge between sample size and estimation error. But the controller design problems have not been considered. Furthermore, [22] considers LQR over an unknown Bernoulli lossy channel, which is estimated using finite samples. A formally equivalent optimal controller is designed using the upper bound of the estimated probability, and its cost is bounded by the maximum estimation error according to the Hoeffding inequality. The setup of [22] is similar to that in this paper.
However, [22] does not explain why the controller relying on the estimated probability stabilizes the system and only provides numerical methods to calculate conservative performance bounds. In fact, if the estimation error is too large, the controller may fail to stabilize the system. Therefore, we seek to gain insight into the LQR problem on the unknown lossy channel. The key issues are determining the maximum estimation error below which the controller stabilizes the system, the minimum number of samples required to ensure stability, and how to analytically express the performance of the estimated controller.
Based on the above analysis, in this paper we consider the LQR problem with unknown Bernoulli packet drop. The control signals may be lost in the controller-actuator channel, resulting in the system input becoming zero at losing instants. The unknown probability is estimated from finite samples, and the estimated probability is used to design an formally equivalent optimal controller. The main contributions are listed below.
  1. 1.
    First, a necessary and sufficient condition is provided for the estimated controller to mean-square stabilize scalar systems, and it is extended to $n$-dimensional systems as a sufficient condition. Based on these conditions, we analyze the stabilizability threshold for the estimation error in $n$-dimensional systems and present its lower bound, below which the estimated controller stabilizes the system. Additionally, we provided tailored results that are less conservative for scalar systems and $n$-dimensional systems with invertible input matrix.
  2. 2.
    We analyze the sample complexity for the stabilizability of the estimated controller, provide an upper bound on the sample complexity, and improve the bounds for scalar systems and $n$-dimensional systems with invertible input matrix. Moreover, we present a sufficient condition for determining whether the estimated controller mean-square stabilizes the system with a certain probability, which can be directly applied in practice, as it does not rely on unknown channel information.
  3. 3.
    We provide an analytical expression for the performance of the estimated controller and the gap between its performance and the optimal performance. Furthermore, we prove that this gap is bounded by the estimation error of the packet drop probability and converges to zero as the estimation error approaches zero.
The remainder of this paper is organized as follows. Section 2 introduces problem formulations; Section 3 analyzes estimation error threshold; Section 4 presents sample complexity; Section 5 provides performance analysis; Section 6 offers numerical simulations; Section 7 concludes the paper.
Notation: $\mathbb{R}^{n\times m}$ is the $n\times m$ dimensional real matrix set; $\mathbb{R}_{+}$ is the set of non-negative real numbers; $\rho(M)$ is the spectral radium of a matrix $M\in\mathbb{R}^{n\times n}$; $\lambda_{\min}(M)$ is the minimum eigenvalue of a real symmetric matrix $M$; $\lambda_{\max}(M)$ is the maximum eigenvalue of a real symmetric matrix $M$; $M^{\top}$ is the transpose of a matrix $M$; $\mathcal{I}$ is the identity matrix with appropriate dimensions; $\mathcal{P}(\cdot)$ is probability measure; $\log(\cdot)$ is the logarithm of a positive real number in base $e$.
2 preliminaries
2.1 Problem Formulation
Considering the linear time-invariant system
$x_{t+1}=Ax_{t}+\lambda_{t}Bu_{t},~{}t=0,1,2,\dots$ (1)
where $x_{t}\in\mathbb{R}^{n}$ is the system state, $x_{0}$ is the initial value with finite expectation and variance, $u_{t}\in\mathbb{R}^{m}$ is the control input, $A\in\mathbb{R}^{n\times n}$ is the state matrix, $B\in\mathbb{R}^{n\times m}$ is the input matrix; $\{\lambda_{t}\}_{t=0}^{+\infty}$ is a sequence of i.i.d. Bernoulli random variables satisfying the following distribution
$\lambda_{t}=\left\{\begin{array}[]{cc}0~{}~{}~{}w.p.&q\\ 1~{}~{}~{}w.p.&1-q\end{array}\right.,$ (2)
where $q\in(0,1)$. The Bernoulli process $\{\lambda_{t}\}_{t=0}^{+\infty}$ models the packet drop phenomenon that may occur in the communication channel [10]. At each instant, the packet of control signal may be lost with a probability of $q$, or successfully transmitted to the system with a probability of $1-q$. According to the definitions of system (1) and (2), if the control packet is lost, the system input becomes zero at the current time. The closed-loop system with packet drop is shown in Fig. 1.
Figure 1: The closed-loop system with packet drop.
The cost function is
$J(x_{0},\textbf{u})=\mathbb{E}\Bigg{\{}\sum_{t=0}^{\infty}x_{t}^{\top}Qx_{t}+\lambda_{t}u_{t}^{\top}Ru_{t}\Bigg{\}},$ (3)
where $\textbf{u}=\{u_{0},u_{1},u_{2},\dots\}$ is the control input sequence, $Q\in\mathbb{R}^{n\times n},R\in\mathbb{R}^{m\times m}$ are positive-definite matrices, and the expectation is taken over the packet drop sequence.
The objective of the LQR problem is to design a control sequence $\hat{\mathbf{u}}$ that mean-square stabilizes the system (1), i.e., $\lim_{t\to\infty}\mathbb{E}[x_{t}^{\top}x_{t}]=0$, and minimizes the cost function $J(x_{0},\textbf{u})$ [25]. We denote the minimum cost function as $J^{*}(x_{0})$ and the optimal control as $\textbf{u}^{*}=\{u_{0}^{*},u_{1}^{*},u_{2}^{*},\dots\}$, i.e., $J^{*}(x_{0})=\min_{\textbf{u}}J(x_{0},\textbf{u})=J(x_{0},\textbf{u}^{*})$. For the infinite horizon LQR problem with Bernoulli packet drop [10, 11], the optimal controller depends on prior knowledge of the true packet drop probability $q$, which is $u^{*}_{t}=Kx_{t}$, where $K=-(R+B^{\top}PB)^{-1}B^{\top}PA$, and $P$ is a $n$-dimensional positive-definite matrix satisfying the modified Riccati equation
$P=Q+A^{\top}PA-(1-q)A^{\top}PB(R+B^{\top}PB)^{-1}B^{\top}PA.$ (4)
And the optimal cost $J^{*}(x_{0})=\mathbb{E}[x_{0}^{\top}Px_{0}]$. To ensure the existence of the positive-definite solution to the modified Riccati equation (4), it is necessary to assume that $(A,B)$ is stabilizable. Moreover, the probability of packet drop cannot exceed a certain critical value [11], which is stated in the following assumption.
Assumption 1
The packet drop probability $q$ satisfies $q<q_{c}$, where $q_{c}$ satisfies
$\frac{1}{\prod_{i}|\lambda_{i}^{u}(A)|^{2}}\leq q_{c}\leq\frac{1}{\max_{i}|\lambda_{i}^{u}(A)|^{2}},$ (5)
and $\lambda_{i}^{u}(A)$ denote the unstable eigenvalues of $A$.
In Assumption 1, $q_{c}$ is an upper bound for $q$ to ensure the existence of a positive-definite solution $P$ to the equation (4), which depends on the system matrix $A$ but cannot be expressed in explicit form. It lies within the certain range indicated in (5). Specially, $q_{c}=1/\max_{i}|\lambda^{u}_{i}(A)|^{2}$ when $B$ is an invertible matrix; $q_{c}=1/\prod_{i}|\lambda_{i}^{u}(A)|^{2}$ when $B$ has rank one [11].
The dependence of $P$ on $q$ indicates that $q$ plays a critical role in obtaining the optimal controller $u^{*}_{t}$. In practice, $q$ may be unknown, and we can only use channel samples to estimate $q$. Let $\hat{q}$ denoted the estimated value of $q$. Given a sample sequence $\{\lambda_{i},i=1,\dots,N_{q}\}$ of i.i.d. Bernoulli random variables, $\hat{q}$ is computed as $\hat{q}=\frac{1}{N_{q}}\sum_{i=1}^{N_{q}}(1-\lambda_{i})$ [22, 26]. Then, we denote the formally equivalent optimal controller designed with $\hat{q}$ as $\hat{\textbf{u}}=\{\hat{u}_{0},\hat{u}_{1},\hat{u}_{2},\dots\}$ and $\hat{u}_{t}$ is
$\hat{u}_{t}=\hat{K}x_{t},~{}t=0,1,2,\dots$ (6)
where
$\hat{K}=-(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A,$ (7)
and $\hat{P}\in\mathbb{R}^{n\times n}$ is the positive-definite solution of the following modified Riccati equation
$\hat{P}=Q+A^{\top}\hat{P}A-(1-\hat{q})A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A.$ (8)
The number of samples $N_{q}$ should ensure $\hat{q}<q_{c}$, so that (8) has a positive-definite solution $\hat{P}$. We will refer to $\hat{K}$ and $\hat{u}$ interchangeably as the equivalent optimal controller.
Since samples are finite, there inevitably exists an estimation error between $q$ and $\hat{q}$. The next subsection will demonstrate that when the error is too large, $\hat{K}$ does not mean-square stabilize the system.
2.2 Motivating Example
In this subsection, we consider the scalar case and provide the necessary and sufficient condition for (7) to mean-square stabilize the system (1) when $n=m=1$. It is straightforward to verify that the necessary and sufficient condition is violated when the estimation error is large.
Theorem 1
When $n=m=1$, with the Assumption 1 and $\hat{q}<q_{c}$ holding, (7) mean-square stabilizes the system (1) if and only if
$Q+(1-q)R\hat{K}^{2}+(\hat{q}-q)(R+B^{2}\hat{P})^{-1}A^{2}B^{2}\hat{P}^{2}>0.$ (9)
Moreover, if $\hat{q}\geq q$, (9) holds.
Proof 2.2.
When $n=m=1$, (7) and (8) can be rewritten as
$\hat{K}=-(R+B^{2}\hat{P})^{-1}AB\hat{P}.$ (10)
$\hat{P}=Q+A^{2}\hat{P}-(1-\hat{q})(R+B^{2}\hat{P})^{-1}A^{2}B^{2}\hat{P}^{2},$ (11)
If the scalar system (1) is mean-square stabilized via (7), we have $\mathbb{E}\{x_{t}^{2}\}\to 0$ as $t\to\infty$, which is equivalent to $\mathbb{E}\{\hat{P}x_{t}^{2}\}\to 0$ as $t\to\infty$, where $\hat{P}$ is a positive real number satisfying (11). Since $\mathbb{E}\{\hat{P}x_{t+1}^{2}\}=[A^{2}+(1-q)(B^{2}\hat{K}^{2}+2AB\hat{K})]^{t+1}\mathbb{E}\{\hat{P}x^{2}_{0}\}$ and $\lim_{t\to\infty}\mathbb{E}\{\hat{P}x_{t}^{2}\}=0$, we have $A^{2}+(1-q)(B^{2}\hat{K}^{2}+2AB\hat{K})<1$ and $\mathbb{E}\{\hat{P}x_{t}^{2}\}$ is monotonically decreasing, which implies that for any $t\geq 0$, $\mathbb{E}\{\hat{P}x_{t+1}^{2}\}<\mathbb{E}\{\hat{P}x_{t}^{2}\}$. On the other hand, we have
$\mathbb{E}\{\hat{P}x_{t+1}^{2}\}$ $=\mathbb{E}\{[A^{2}+(1-q)(B^{2}\hat{K}^{2}+2AB\hat{K})]\hat{P}x_{t}^{2}\}$
=$\overset{(a)}{=}\mathbb{E}\{\hat{P}x_{t}^{2}\}-[Q+(1-q)R\hat{K}^{2}+(\hat{q}-q)\times$
$\quad(R+B^{2}\hat{P})^{-1}A^{2}B^{2}\hat{P}^{2}]\mathbb{E}\{x_{t}^{2}\},$ (12)
where $(a)$ follows from (10),(11). Therefore, $\mathbb{E}\{\hat{P}x_{t+1}^{2}\}<\mathbb{E}\{\hat{P}x_{t}^{2}\}$ is equivalent to (9). Moreover, if $\hat{q}\geq q$, (9) always holds because all terms are positive.
Theorem 1 implies that $\hat{K}$ mean-square stabilizes the scalar system (1) when $q-\hat{q}\leq 0$. In fact, this conclusion also holds for $n$-dimensional systems, which will be proven later. However, when $q-\hat{q}>0$, it is easy to verify that (9) may be violated. For example, $q=0.4$, $\hat{q}=0$, $A=1.5$ and $B=Q=R=1$. Therefore, when the estimated error $q-\hat{q}$ is large, the controller (7) designed with $\hat{q}$ does not mean-square stabilize the system (1). Inspired by the above example, this paper focuses on three problems: for $\hat{K}$ stabilizing the system, what is the upper bound of the estimation error $q-\hat{q}$, what is the minimum sample size required, and how to assess the performance of $\hat{K}$. The answers to the three questions above will be provided in Sections 3, 4, and 5, respectively.
3 Stabilizability analysis
In this section, we first analyze the stabilizability threshold of $q-\hat{q}$ for general $n$-dimensional systems, below which $\hat{K}$ mean-square stabilizes the system. Then, tailored results with less conservatism are given for some specific cases.
3.1 General results
First, we provide the following theorem which guarantees (7) mean-square stabilizes the system (1).
Theorem 3.3.
With Assumption 1 and $\hat{q}<q_{c}$ holding, (7) mean-square stabilizes the system (1), if $q,\hat{q}$ satisfy
$Q+(1-q)\hat{K}^{\top}R\hat{K}-(q-\hat{q})A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A>0$ (13)
where $\hat{P}$ is the positive-definite solution of (8). Moreover, if $\hat{q}{\geq}q$, (13) holds.
Proof 3.4.
Introducing a Lyapunov function $\hat{V}(x_{t})=\mathbb{E}[x_{t}^{\top}\hat{P}x_{t}]$, and we can derive that
$\hat{V}(x_{t+1})=$ $\mathbb{E}\Big{\{}(Ax_{t}+\lambda_{t}B\hat{u}_{t})^{\top}\hat{P}(Ax_{t}+\lambda_{t}B\hat{u}_{t})\Big{\}}$
=$\overset{(a)}{=}$ $\mathbb{E}\Big{\{}x_{t}^{\top}\big{[}A^{\top}\hat{P}A-(1-q)\hat{K}^{\top}R\hat{K}-(1-q)\times$
$A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A\big{]}x_{t}\Big{\}}$
=$\overset{(b)}{=}$ $\mathbb{E}\Big{\{}x_{t}^{\top}\big{[}\hat{P}-Q-(1-q)\hat{K}^{\top}R\hat{K}+(q-\hat{q})\times$
$A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A\big{]}x_{t}\Big{\}}$
<$\overset{(c)}{<}$ $\mathbb{E}\{x_{t}^{\top}\hat{P}x_{t}\}=\hat{V}(x_{t}).$ (14)
where $(a)$, $(b)$, $(c)$ follow from (6), (8), (13), respectively.
Based on Lyapunov theory, we derive $\lim_{t\to\infty}\mathbb{E}\{x^{\top}_{t}x_{t}\}=0$, thus proving the mean-square stability of the system (1). Moreover, since $Q>0$, $(1-q)\hat{K}^{\top}R\hat{K}>0$, $A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A\geq 0$, (13) always holds if $\hat{q}{\geq}q$.
The condition (13) is a generalization of (9) from scalar systems to $n$-dimensional systems. It is worth noting that (9) is a necessary and sufficient condition for $\hat{K}$ to stabilize the system in the scalar case. However, for $n$-dimensional systems, (13) is only a sufficient condition, as $\mathbb{E}\{x^{\top}_{t+1}x_{t+1}\}<\mathbb{E}\{x^{\top}_{t}x_{t}\}$ is not equivalent to $\mathbb{E}\{x^{\top}_{t+1}\hat{P}x_{t+1}\}<\mathbb{E}\{x^{\top}_{t}\hat{P}x_{t}\}$ for the given matrix $\hat{P}$.
Since a larger $q$ represents a more hostile communication environment, where more packets may be lost, the second conclusion of Theorem 3.3 indicates that the system (1) can be stabilized in a more conservative manner. In other words, it can be stabilized via the controller (7) designed with a $\hat{q}$ greater than the true probability. But this conservative manner may incur a higher cost, which will be validated in the simulations in Section 5, as shown in Fig. 2(b). Then, we focus on the stabilizability threshold for the estimation error $\hat{q}<q$.
For the sake of clarity, we define the Stabilizability Threshold (ST) as
$\bar{\delta}=\max\Big{\{}\delta\in\mathbb{R}_{+}\Big{|}\mathcal{C}(q,\hat{q})>0,~{}\forall(q,\hat{q})\in\Omega_{\delta}\Big{\}}.$ (15)
where $\mathcal{C}(q,\hat{q})=Q+(1-q)\hat{K}^{\top}R\hat{K}-(q-\hat{q})A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A$, i.e., the left-hand side of (13), and $\Omega_{\delta}=\big{\{}(q,\hat{q})\in\mathbb{R}\times\mathbb{R}\big{|}q-\hat{q}<\delta,~{}0\leq q,\hat{q}<q_{c}\big{\}}$.
From the definition (15), we know that when $q-\hat{q}<\bar{\delta}$, $\hat{K}$ mean-square stabilizes the system according to the Theorem 3.3. We aim to determine the ST for a given LQR problem (1)(2)(3), i.e., given $A,B,Q,R,q$. Because ST may depends on all these parameters, it is appropriate to denote it as $\bar{\delta}(A,B,Q,R,q)$. And we abbreviate it as $\bar{\delta}(q)$, since $q$ is the essential cause of the estimation error.
In fact, it is challenging to analytically express the ST $\bar{\delta}(q)$ in terms of the given parameters $A,B,Q,R$, and $q$, even in the scalar case. Therefore, in this section, we will provide a lower bound for the ST. This lower bound can be expressed explicitly based on the given parameters and ensures that (7) mean-square stabilizes the system (1) when $q-\hat{q}$ is smaller than this lower bound. We first show a useful lemma.
Lemma 3.5.
With Assumption 1 and $\hat{q}<q_{c}$ holding, given $\hat{q}<q$, then the corresponding matrices $P$ and $\hat{P}$ in (4) and (8) satisfy $\hat{P}\leq P$.
Proof 3.6.
Let’s define $g(X)\triangleq Q+A^{\top}XA-(1-q)A^{\top}XB(R+B^{\top}XB)^{-1}B^{\top}XA$, $\hat{g}(X)\triangleq Q+A^{\top}XA-(1-\hat{q})A^{\top}XB(R+B^{\top}XB)^{-1}B^{\top}XA$, where $X\in\mathbb{R}^{n\times n}$. Given any matrix $\mathcal{P}_{0}>0$ as the initial matrix and denoting $P_{t+1}=g(P_{t}),~{}\hat{P}_{t+1}=g(\hat{P}_{t})$, we know from [9] that $P=\lim_{t\to\infty}P_{t}=\lim_{t\to\infty}g^{t}(\mathcal{P}_{0})$, $\hat{P}=\lim_{t\to\infty}\hat{P}_{t}=\lim_{t\to\infty}\hat{g}^{t}(\mathcal{P}_{0})$.
When $\hat{q}<q$, i.e. $1-\hat{q}>1-q$, we know from c) of Lemma 1 in [9] that $\hat{P}_{1}=\hat{g}(\mathcal{P}_{0})\leq g(\mathcal{P}_{0})=P_{1}$. And from d) of Lemma 1 in [9], we have $\hat{g}(\hat{P}_{1})\leq g(\hat{P}_{1})\leq g(P_{1})$, which shows $\hat{P}_{2}\leq P_{2}$. By induction, there is $\hat{P}_{t}\leq P_{t}$ for $\forall t$. Let $t\to\infty$, we have $\hat{P}\leq P$.
Then, the lower bound of ST is given as follows.
Theorem 3.7.
With Assumption 1 and $\hat{q}<q_{c}$ holding, for the given system (1), the ST satisfies
$\bar{\delta}(q)\geq\frac{\lambda_{\min}\{Q^{\frac{1}{2}}(A^{\top}P^{2}A)^{-1}Q^{\frac{1}{2}}\}}{c_{1}},$ (16)
where $c_{1}=\frac{\lambda_{\max}(BB^{\top})}{\lambda_{\min}(R+B^{\top}P_{0}B)}$, $P$ is the positive-definite solution of (4), $P_{0}$ is the positive-definite solution of the following Riccati equation
$P_{0}=Q+A^{\top}P_{0}A-A^{\top}P_{0}B(R+B^{\top}P_{0}B)^{-1}B^{\top}P_{0}A.$ (17)
Proof 3.8.
For any $q$ and $\hat{q}$ satisfying $0<q-\hat{q}<\lambda_{\min}\{Q^{\frac{1}{2}}(A^{\top}P^{2}A)^{-1}Q^{\frac{1}{2}}\}/{c_{1}}$, we know $P_{0}\leq\hat{P}\leq P$ from Lemma 3.5 and $Q-(q-\hat{q})c_{1}A^{\top}P^{2}A>0$, where $\hat{P}$ satisfies (8). Then, we have
$Q\!+\!(1\!-\!q)\hat{K}^{\top}\!R\hat{K}\!-\!(q\!-\!\hat{q})A^{\top}\!\hat{P}B(\!R\!+\!B^{\top}\!\hat{P}\!B)^{-1}\!B^{\top}\!\hat{P}\!A$ (18)
$\geq$ $Q-(q-\hat{q})\frac{\lambda_{\max}(BB^{\top})}{\lambda_{\min}(R+B^{\top}P_{0}B)}A^{\top}\hat{P}^{2}A$
$\geq$ $Q-(q-\hat{q})c_{1}A^{\top}P^{2}A>0.$
According to Theorem 3.3, (7) mean-square stabilizes (1).
Remark 3.9.
If the true probability $q$ increases, the upper bound on $q-\hat{q}$ becomes more restrictive due to the increase of $P$ with respect to $q$. In particular, if $q$ approaches $q_{c}$, the upper bound will approach 00. Therefore, in poorer communication channels, the estimate should be more accurate to ensure $\hat{K}$ mean-square stabilizes the system. As shown in Fig. 3(b),4(b), the closer $q$ is to $q_{c}$, the smaller the range of stabilizable $\hat{q}$.
From Theorem 3.7, we know that when $q$ is smaller than the lower bound of the ST, (7) designed with any $\hat{q}\in[0,q_{c})$ mean-square stabilizes the system (1), as shown below.
Corollary 3.10.
With Assumption 1 and $\hat{q}<q_{c}$ holding, for any $\hat{q}\in[0,q_{c})$, (7) mean-square stabilizes the system (1) if $q$ satisfies $q<\lambda_{\min}\{Q^{\frac{1}{2}}(A^{\top}P^{2}A)^{-1}Q^{\frac{1}{2}}\}/{c_{1}}$, where the definitions of $c_{1},P,P_{0}$ are the same as those in Theorem 3.7.
Theorem 3.7 provides a lower bound for the ST for general $n$-dimensional systems. Moreover, tighter lower bounds can be obtained for scalar systems and $n$-dimensional systems with invertible input matrix.
3.2 Tailored results for specific classes of systems
First, the lower bound of ST is given for scalar systems.
Theorem 3.11.
With Assumption 1 and $\hat{q}<q_{c}$ holding, for the given system (1), supposing $n=m=1$, the ST satisfies
$\bar{\delta}(q)\geq\frac{Q(R+B^{2}P)}{A^{2}B^{2}P^{2}}+\frac{(1-q)R}{R+B^{2}P},$ (19)
where $P$ is the positive-definite solution of (4).
Proof 3.12.
Through algebraic manipulation, we derive that (9) in Theorem 1 is equivalent to $q-\hat{q}<\frac{Q(R+B^{2}\hat{P})}{A^{2}B^{2}\hat{P}^{2}}+\frac{(1-q)R}{R+B^{2}\hat{P}}$. Define $f(X)\triangleq\frac{Q(R+B^{2}X)}{A^{2}B^{2}X^{2}}+\frac{(1-q)R}{R+B^{2}X}$ where $X\in\mathbb{R}_{+}$. Since $Q,R>0$, $f(X)$ is monotonically decreasing with respect to $X\in\mathbb{R}_{+}$. From Lemma 3.5, we know $f(\hat{P})\geq f(P)$ if $\hat{q}<q$. Hence, when $0\leq q-\hat{q}\leq\frac{Q(R+B^{2}P)}{A^{2}B^{2}P^{2}}+\frac{(1-q)R}{R+B^{2}P}$, (9) holds.
Remark 3.13.
In the scalar case, Theorem 3.11 is less conservative than Theorem 3.7. In fact, for scalar case, the right-hand side (RHS) of (16) can be rewritten as $\frac{Q(R+B^{2}P_{0})}{A^{2}B^{2}P^{2}}$ which is clearly smaller than the RHS of (19). Moreover, from Theorem 3.11, when $q<\frac{Q(R+B^{2}P)}{A^{2}B^{2}P^{2}}+\frac{(1-q)R}{R+B^{2}P}$, (7) designed with any $\hat{q}\in[0,1/\rho(A)^{2})$ mean-square stabilizes the system (1).
Next, the lower bound of ST is presented for $n$-dimensional systems with invertible $B$.
Theorem 3.14.
With Assumption 1 and $\hat{q}<q$ holding, for the given system (1), supposing $B$ is invertible, the ST satisfies
$\bar{\delta}(q)\geq\lambda_{\min}\{\Xi(A^{\top}PA)^{-1}\Xi\},$ (20)
where $\Xi=(Q+(1-q)c_{2}A^{\top}P_{0}^{2}A)^{\frac{1}{2}}$, $c_{2}=\frac{\lambda_{\min}(R)\lambda_{\min}(BB^{\top})}{\lambda_{\max}(R+B^{\top}PB)^{2}}$, $P$, $P_{0}$ are defined as in Theorem 3.7.
Proof 3.15.
When $0<q-\hat{q}<\lambda_{\min}\{\Xi(A^{\top}PA)^{-1}\Xi\}$, we have
$Q\!+\!(1\!-\!q)\hat{K}^{\top}\!R\hat{K}\!-\!(q\!-\!\hat{q})A^{\top}\hat{P}B(R\!+\!B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A$
$\geq$ $Q+(1-q)\frac{\lambda_{\min}(R)\lambda_{\min}(BB^{\top})}{\lambda_{\max}(R+BPB^{\top})^{2}}A^{\top}\hat{P}^{2}A-(q-\hat{q})\times$
$A^{\top}\hat{P}[(BR^{-1}B^{\top})^{-1}+\hat{P}]^{-1}\hat{P}A$
$\overset{(a)}{\geq}$ $Q+(1-q)c_{2}A^{\top}P_{0}^{2}A-(q-\hat{q})A^{\top}PA>0,$ (21)
where $(a)$ is derived from $\hat{P}^{\frac{1}{2}}[(BR^{-1}B^{\top}+\hat{P})]^{-1}\hat{P}^{\frac{1}{2}}\leq\mathcal{I}$ and Lemma 3.5. From Theorem 3.3, (7) mean-square stabilizes (1).
Remark 3.16.
The characteristics in Remark 3.9 still hold for the tailored results, for similar reasons. In a poorer communication channel, as $q$ increases, the upper bounds on $q-\hat{q}$ in Theorem 3.11, 3.14 become more restrictive, as shown in Fig. 3(a) and 4(a).
Remark 3.17.
For systems with invertible $B$, Theorem 3.14 is less conservative than Theorem 3.7. If $q\!-\!\hat{q}\!<\!\lambda_{\min}\!\{\!Q^{\frac{1}{2}}\!(\!A^{\top}\!P^{2}\!A)^{-1}Q^{\frac{1}{2}}\!\}/{c_{1}}$, then $c_{1}(q-\hat{q})A^{\top}P^{2}A<Q$. Since $A^{\top}\!P\!A\!<\!A^{\top}\!P\!B(R\!+\!B^{\top}\!P\!B)^{-1}\!B^{\top}\!P\!A\!\leq\!c_{1}A^{\top}\!P^{2}\!A$ when $B$ is invertible, we know $(q-\hat{q})A^{\top}PA<Q+(1-q)c_{2}A^{\top}P_{0}^{2}A$. So $q-\hat{q}<\lambda_{\min}\{\Xi(A^{\top}PA)^{-1}\Xi\}$. Additionally, Theorem 3.14 also applies to scalar systems, but it is more conservative than Theorem 3.11 in scalar case. It can be verified that, when $n=m=1$, $\lambda_{\min}\{\Xi(A^{\top}PA)^{-1}\Xi\}=\frac{Q+(1-q)RA^{2}B^{2}P_{0}^{2}(R+B^{2}P)^{-2}}{A^{2}P}\leq\frac{Q(B^{-2}R+P)+(1-q)RA^{2}P^{2}(R+B^{2}P)^{-1}}{A^{2}P^{2}}=\frac{Q(R+B^{2}P)}{A^{2}B^{2}P^{2}}+\frac{(1-q)R}{R+B^{2}P}$, where $\Xi$, $P_{0}$ are defined as in Theorem 3.14 and 3.11, respectively. Moreover, from Theorem 3.14, when $q<\lambda_{\min}\{\Xi(A^{\top}PA)^{-1}\Xi\}$, (7) designed with any $\hat{q}\in[0,\frac{1}{\rho(A)^{2}})$ mean-square stabilizes the system (1).
In this section, we present upper bounds of $q-\hat{q}$ for $\hat{K}$ stabilizing the system (1). Essentially, the presence of non-zero estimation error stems from the fact that only finite samples are utilized for estimating the packet drop probability. Next, we will investigate the sample complexity for estimating the packet drop probability such that $\hat{K}$ stabilizes the system (1).
4 samples complexity analysis
According to the strong law of large numbers (Theorem 5.18 of [27]), as $N_{q}$ tends to +$\infty$+ ∞, $\hat{q}$ converges to $q$ almost surely. However, in practice, samples are always finite. Therefore, we are interested in the sample complexity for estimating the packet drop probability. The following lemma is a variation of the Hoeffding’s inequality, which establishes a bridge between sample size and estimation error.
Lemma 4.18.
(Theorem 4.5 of [27]) Given a sample sequence $\{\lambda_{i},i=1,\dots,N_{q}\}$ and letting $\hat{q}=\frac{1}{N_{q}}\sum_{i=1}^{N_{q}}(1-\lambda_{i})$. Then for any $\beta\in(0,1)$, it holds that $\mathcal{P}\{|q-\hat{q}|\leq\Delta(N_{q},\beta)\}\geq 1-\beta$, where $\Delta(N_{q},\beta):=\sqrt{\frac{\log(2/\beta)}{2N_{q}}}$.
In fact, for a certain $N_{q}$, if the upper bound on the estimation error is smaller than the ST, (7) mean-square stabilizes the system (1). Based on this idea, we present the following lemma regarding the sample complexity.
Lemma 4.19.
Under Assumption 1, given a sequence of samples $\{\lambda_{i},i=1,\dots,N_{q}\}$ such that $\hat{q}<q_{c}$, (7) mean-square stabilizes the system (1) with probability at least $1-\beta,~{}\forall\beta\in(0,1)$, if $N_{q}$ satisfies
$N_{q}>\frac{\log(2/\beta)}{2\bar{\delta}(q)^{2}},$ (22)
where $\bar{\delta}(q)$ is the ST for the given system (1).
Proof 4.20.
According to Lemma 4.18, we know that $\mathcal{P}\{-\Delta(N_{q},\beta){\leq}q-\hat{q}\leq\Delta(N_{q},\beta)\}\geq 1-\beta$. Then, based on (22), we have $\Delta(N_{q},\beta)<\bar{\delta}(q)$. Therefore, $q-\hat{q}<\bar{\delta}(q)$ holds with probability at least $1-\beta$.
The RHS of (22) represents the sample complexity for $\hat{K}$ to mean-square stabilize the system (1). Evidently, for a given system, as $\bar{\delta}(q)$ increases, the sample complexity decreases. By integrating the analysis of $\bar{\delta}(q)$ from the previous section, we derive upper bounds on the sample complexity from the system parameters $A,B,Q,R,q$. Next theorem gives an upper bound on the sample complexity for $n$-dimensional systems.
Theorem 4.21.
Under Assumption 1, given a sequence of samples $\{\lambda_{i},i=1,\dots,N_{q}\}$ such that $\hat{q}<q_{c}$, (7) mean-square stabilizes the system (1) with probability at least $1-\beta$, if $N_{q}$ satisfies
$N_{q}>\frac{c_{1}^{2}\log(2/\beta)}{2\lambda_{\min}\{Q^{\frac{1}{2}}(A^{\top}P^{2}A)^{-1}Q^{\frac{1}{2}}\}^{2}},$ (23)
where $\beta\in(0,1)$, and $c_{1},P$ are defined as in Theorem 3.7.
Substituting the lower bound of $\bar{\delta}(q)$ from Theorem 3.7 into (22), Theorem 4.21 can be directly obtained. The RHS of (23) serves as an upper bound of the sample complexity. Furthermore, when $q$ is sufficiently small, the sample size required for stability may be less than (23). Based on Corollary 3.10, when $q\!<\!{\lambda_{\min}\{Q^{\frac{1}{2}}(A^{\top}\!P^{2}\!A)^{-1}Q^{\frac{1}{2}}\}}/{c_{1}}$, $\hat{K}$ stabilizes the system even if $N_{q}=0$. However, when $q$ is large and approaches $q_{c}$, the sample complexity will increase rapidly and tend toward +$\infty$+ ∞, as $P$ increases unboundedly [9].
Additionally, under the same assumptions as in Theorem 4.21, for scalar systems, $\hat{K}$ mean-square stabilizes the system with probability at least $1-\beta$ if $N_{q}$ satisfies
$N_{q}>\frac{\log(2/\beta)A^{4}B^{4}P^{4}(R+B^{2}P)^{2}}{[Q(R+B^{2}P)^{2}+(1-q)RA^{2}B^{2}P^{2}]^{2}};$ (24)
for systems with invertible $B$, $\hat{K}$ stabilizes the system with probability at least $1-\beta$ if $N_{q}$ satisfies
$N_{q}>\frac{\log(2/\beta)}{\lambda_{\min}\{\Xi(A^{\top}PA)^{-1}\Xi\}^{2}}.$ (25)
The results (24) and (25) can be proven by combining Lemma 4.19 with Theorem 3.11 and 3.14, respectively.
Since ST and sample complexity depend on the unknown $q$, they cannot be directly used to determine whether $\hat{K}$ mean-square stabilizes the system in practice. Therefore, from a practical perspective, a sufficient condition is provided, which is independent of the unknown $q$, to assert that $\hat{K}$ means-square stabilizes the system (1) with a certain probability.
Theorem 4.22.
Under Assumption 1, given a sequence of samples $\{\lambda_{i},i\!=\!1,\dots,N_{q}\}$ such that $\hat{q}<q_{c}$, (7) mean-square stabilizes system (1) with probability at least $1-\beta$, if
$\bar{q}\geq\hat{q}+\Delta(N_{q},\beta),$ (26)
where $\beta\in(0,1)$, $\Delta(N_{q},\beta)$ is defined in Lemma 4.18, $\bar{q}$ is the optimal value of the semi-definite programming (SDP)
$\max_{x\in[0,1]}\quad$ $x$ (27a)
${\rm s.t.}\quad$ $x\geq\hat{q}$ (27b)
$Q+(1-x)\hat{K}^{\top}R\hat{K}-(x-\hat{q})\times$
$A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A>0,$ (27c)
where $\hat{K}$, $\hat{P}$ are defined in (7), (8), respectively.
Proof 4.23.
According to Lemma 4.18, we have $\mathcal{P}\{-\Delta(N_{q},\beta){\leq}q-\hat{q}\leq\Delta(N_{q},\beta)\}\geq 1-\beta.$ Combining with (26), we know $q\leq\bar{q}$ with probability at least $1-\beta$. Hence, there is $Q\!+\!(1\!-\!q)\hat{K}^{\top}\!R\hat{K}\!-\!(q\!-\!\hat{q})A^{\top}\!\hat{P}B(R\!+\!B^{\top}\!\hat{P}\!B)^{-1}\!B^{\top}\!\hat{P}\!A>\!Q\!+\!(1\!-\!\bar{q})\hat{K}^{\top}\!R\hat{K}\!-\!(\bar{q}-\hat{q})A^{\top}\!\hat{P}B(R\!+\!B^{\top}\!\hat{P}\!B)^{-1}B^{\top}\!\hat{P}\!A$ with probability at least $1-\beta$. Then, from the constraint (27c), we know that $Q+(1-q)\hat{K}^{\top}R\hat{K}-(q-\hat{q})A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A>0$ with probability at least $1-\beta$. According to Theorem 3.3, (7) mean-square stabilizes the system (1) with probability at least $1-\beta$.
Remark 4.24.
Theorem 4.22 provides a sufficient condition for $\hat{K}$ mean-square stabilizing the system (1). Given a $\hat{q}$ estimated from $N_{q}$ samples, if (26) holds, then $\hat{K}$ mean-square stabilizes the system with probability at least $1-\beta$. Since (26) is independent of the unknown $q$, it can be directly applied in practice. It is important to emphasize that $\bar{q}$ is only used to verify the validity of (26) and is not involved in constructing the controller. Therefore, even a large $\bar{q}$ does not introduce additional conservatism into this theorem. Moreover, the SDP (27) is always feasible. For instance, $x=\hat{q}$ is a feasible solution as it clearly satisfies the constraints (27b) and (27c).
A numerical example is provided to illustrate the applicability of Theorem 4.22. Consider the two-dimensional system $(A,B)$ given in Example 6.31, initial value $x_{0}=[0.9325;1.1616]$. Assuming $q=0.2$ which satisfies Assumption 1 and is unknown to the controller. Using $300$ samples, we obtain an estimated probability $\hat{q}=0.1633$. Substituting $\hat{q}=0.1633$ into (27b) and (27c), the solution to (27) is $\bar{q}=0.4181$. It can be verified that (26) holds when $\beta=0.01$, and then we know $\hat{K}$ designed with $\hat{q}=0.1633$ mean-square stabilizes the system with probability at least $0.99$. The system trajectory of Example 6.31 under $\hat{K}$ is displayed in Fig. 2(a).
(a) Given $(A,B)$ as in Example 6.31 and $q=0.2$, $\hat{K}$ designed with $\hat{q}=0.1633$ stabilizes the system, which is estimated by $300$ samples.
(b) For $(A,B)$ in Example 6.31, the gap between the cost of $\hat{K}$ and the optimal cost, i.e., $J(x_{0},\hat{\mathbf{u}})$ and $J^{*}(x_{0})$.
Figure 2: The trajectory and gap between $J(x_{0},\hat{\mathbf{u}})$ and $J^{*}(x_{0})$ for the system given in Example 6.31 with $q=0.2$.
5 Performance analysis
The stability issue under unknown packet drop probability has been discussed earlier. Next, the performance of $\hat{K}$ is also an important concern. In this section, we analytically express the performance of $\hat{K}$ and bound the gap between its performance and the optimal performance $J^{*}(x_{0})$. Then, prove that the gap converges to 00 as $\hat{q}$ approaches to $q$. First, the continuity of $\hat{P}$ with respect to $\hat{q}$ is given in next lemma.
Lemma 5.25.
With Assumption 1 and $\hat{q}<q_{c}$ holding, if $\hat{q}$ converges to $q$, then the corresponding matrix $\hat{P}$ converges to $P$, i.e., $\lim_{\hat{q}\to q}\hat{P}=P$, where $\hat{P}$, $P$ are the positive-definite solutions of (4) and (8) respectively.
Proof 5.26.
When $\hat{q}<q$ and $\hat{q}$ increases towards $q$, according to Lemma 3.5, the corresponding matrix $\hat{P}$ is monotonically increasing and bounded. Therefore, there exists a positive-definite matrix $\hat{P}^{-}$ such that $\lim_{\hat{q}{\to}q^{-}}\hat{P}=\hat{P}^{-}$. Then, taking the left limit with respect to $\hat{q}$ on both side of (8), we have $\hat{P}^{-}\!=\!Q\!+\!A^{\top}\!\hat{P}^{-}\!\!A\!-\!(1\!-q)A^{\top}\!\hat{P}^{-}\!B(\!R\!+\!B^{\top}\!\hat{P}^{-}\!B)^{-1}\!B^{\top}\!\hat{P}^{-}\!\!A$ which shows $\hat{P}^{-}$ is the positive solution of (4). We have $\hat{P}^{-}=P$ due to the uniqueness of the solution of (4). Similarly, when $\hat{q}>q$ and $\hat{q}$ decreases towards $q$, there also exists a positive-definite matrix $\hat{P}^{+}$ satisfying $\lim_{\hat{q}{\to}q^{+}}\hat{P}=\hat{P}^{+}$ and it can be demonstrated that $\hat{P}^{+}=P$. In summary, $\lim_{\hat{q}{\to}q}\hat{P}=P$.
Then, next theorem analytically expresses the gap between the cost of $\hat{K}$ and the optimal cost $J^{*}(x_{0})$.
Theorem 5.27.
With Assumption 1 and $\hat{q}<q_{c}$ holding, when (7) mean-square stabilizes the system (1), the gap between $J(x_{0},\mathbf{\hat{u}})$ and the optimal cost $J^{*}(x_{0})$ satisfies
$J(x_{0},\hat{\mathbf{u}})-J^{*}(x_{0})=$ ${\rm tr}\big{\{}(q-\hat{q})X_{\hat{K}}+(\hat{P}-P)X_{0}\big{\}}$ (28)
where $X_{\hat{K}}=A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}A\mathbb{E}\{\sum_{t=0}^{\infty}x_{t}x_{t}^{\top}\}$, $X_{0}=\mathbb{E}(x_{0}x_{0}^{\top})$ are bounded positive semi-definite matrices. Moreover, as $\hat{q}$ converges to $q$, the gap $J(x_{0},\mathbf{\hat{u}})-J^{*}(x_{0})$ converges to 00.
Proof 5.28.
When (7) stabilizes the system (1), the cost $J(x_{0},\mathbf{\hat{u}})=\sum_{t=0}^{\infty}\mathbb{E}\Big{\{}x_{t}^{\top}Qx_{t}+\lambda_{t}\hat{u}_{t}R\hat{u}_{t}\Big{\}}$ is finite, and $\lim_{t\to\infty}\mathbb{E}\{x_{t+1}^{\top}\hat{P}x_{t+1}\}=0$. Therefore,
$J(x_{0},\mathbf{\hat{u}})=\lim_{t\to\infty}\sum_{n=0}^{t}\mathbb{E}\{x_{n}^{\top}Qx_{n}+\lambda_{n}\hat{u}_{n}^{\top}R\hat{u}_{n}+x_{t+1}^{\top}\hat{P}x_{t+1}\}$
=$\overset{(a)}{=}$ $\lim_{t\to\infty}\Bigg{[}\sum_{n=0}^{t-1}\mathbb{E}\{x_{n}^{\top}Qx_{n}+\lambda_{n}\hat{u}_{n}^{\top}R\hat{u}_{n}+x_{t}^{\top}\hat{P}x_{t}\}$
$+(q-\hat{q})\mathbb{E}\{x_{t}^{\top}A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}Ax_{t}\}\Bigg{]}$
=$\overset{(b)}{=}$ $\lim_{t\to\infty}\Bigg{[}(q-\hat{q})\sum_{n=0}^{t}\mathbb{E}\{x_{n}^{\top}A^{\top}\hat{P}B(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}Ax_{n}\}$
$+$ $\mathbb{E}\{x_{0}^{\top}\hat{P}x_{0}\}\Bigg{]}={\rm tr}\big{\{}(q-\hat{q})X_{\hat{K}}+\hat{P}X_{0}\big{\}},$ (29)
where $(a)$ follows from $\hat{u}_{t}=-(R+B^{\top}\hat{P}B)^{-1}B^{\top}\hat{P}Ax_{t}$; $(b)$ follows from substituting $\hat{u}_{t-1},\hat{u}_{t-2},\dots,\hat{u}_{0}$ in sequence. Because $\hat{K}$ mean-square stabilizes the system (1), $\mathbb{E}\{\sum_{t=0}^{\infty}x_{t}x_{t}^{\top}\}$ is bounded. Based on (5.28) and $J^{*}(x_{0})={\rm tr}\{PX_{0}\}$, we obtain (28). Moreover, since $X_{\hat{K}},X_{0}$ are bounded and $\lim_{\hat{q}{\to}q}\hat{P}=P$ (given by Lemma 5.25), there is $\lim_{\hat{q}\to q}J(x_{0},\hat{\mathbf{u}})-J^{*}(x_{0})=0$.
(a) Stabilizability region according to Theorem 3.11.
(b) Stabilizability region according to Theorem 3.7.
(c) Sample complexity based on (24).
(d) Sample complexity based on (23).
Figure 3: Given the scalar system of Example 6.30. In Fig. 3(a) and 3(b), blue region represents $\hat{K}$ designed with $\hat{q}$ in it stabilizes the system under the corresponding $q$; red region indicates $\hat{K}$ cannot stabilize the system; gray region is not classifiable by Theorem 3.7,3.11. If $\beta=0.1$, Fig. 3(c),3(d) represent the sample complexity for the scalar system in Example 6.30. From (24), when $q<0.231$, $\hat{K}$ designed using any $\hat{q}\in[0,1/\rho(A)^{2})$ stabilizes the system. And (23) implies when $q<0.128$, $\hat{K}$ with any $\hat{q}\in[0,1/\rho(A)^{2})$ stabilizes the system.
Remark 5.29.
Based on Theorem 5.27, more intuitive and concise upper bounds on the cost gap can be derived from (28). Notice that, when $\hat{q}<q$, we have $J(x_{0},\hat{\mathbf{u}})-J^{*}(x_{0})\leq{\rm tr}\{X_{\hat{K}}\}(q-\hat{q})$ because $\hat{P}-P\leq 0$ (proven by Lemma 1). And when $\hat{q}>q$, there is $J(x_{0},\hat{\mathbf{u}})-J^{*}(x_{0})\leq{\rm tr}\{X_{0}\}\lambda_{\max}\{\hat{P}-P\}$. Therefore, the cost gap can always be directly bounded by the estimated error $q-\hat{q}$ or by the corresponding matrices $\hat{P}-P$. Consider the system given in Example 6.31 with $x_{0}\!=\![5;5]$ and assume the true packet probability $q=0.2$. Fig. 2(b) illustrates the gap $J(x_{0},\hat{\mathbf{u}})-J^{*}(x_{0})$ under different estimated probability $\hat{q}$. As $\hat{q}$ approaches $q$, the gap decreases, whereas as $\hat{q}$ diverges from $q$, the gap increases. According to the above analysis, if the sample size approaches infinity, the cost gap tends to zero almost surely. Furthermore, as shown in Fig. 2(b), using a larger estimate to design the controller is a conservative approach that may incur higher costs. For examole, the cost at $\hat{q}=0.4$ is clearly greater than the cost at $\hat{q}=0$.
(a) Stabilizability region according to Theorem 3.14.
(b) Stabilizability region according to Theorem 3.7.
(c) Sample complexity based on (25).
(d) Sample complexity based on (23).
Figure 4: Given the two-dimensional system of Example 6.31. In Fig. 4(a), 4(b), the blue, red and gray regions have the same meaning for Theorem 3.7, 3.14 as they do for Theorem 3.7,3.11 in previous example. Fig. 4(c), 4(d) illustrate sample complexity when $\beta=0.1$. From (25), when $q<0.167$, $\hat{K}$ with any $\hat{q}\in[0,1/\rho(A)^{2})$ stabilizes the system. From (23), when $q<0.104$, $\hat{K}$ with any $\hat{q}\in[0,1/\rho(A)^{2})$ stabilizes the system.
6 Simulations
This section presents numerical simulations of two systems, illustrating the stabilizing range of $\hat{q}$ under different $q$ and sample complexity for stabilizability.
Example 6.30.
Consider a scalar system with $A=1.5$, $B=Q=R=1$. The stabilizability regions and sample complexities are shown in Fig. 3. Based on the necessary and sufficient condition in Theorem 1, the unstabilizable regions (red regions) in Fig. 3(a),3(b) are searched with a step size $0.001$ of $q$ and $\hat{q}$. The stabilizable regions (blue regions) in Fig. 3(a),3(b) are determined by Theorem 3.11 and 3.7, respectively. The error bounds provided by Theorem 3.7,3.11 for scalar systems under different $q$ are presented in Fig. 5. The gray area should ideally be blue, but it cannot be classified by the Theorem 3.11 and 3.7. The conservatism curves of Theorem 3.11 and 3.7 is depicted in Fig. 6. For scalar systems, Theorem 3.11 exhibits less conservatism than Theorem 3.7, as Theorem 3.11 is the tailored result. Hence, the sample complexity based on the tailored result (24) is also significantly lower than that based on the general result 23, as shown in Fig. 3(c),3(d).
Example 6.31.
Given a two-dimensional system with $A=[1.5,0.1;0,1]$, $B=Q=R=[1,0;0,1]$ According to [28], for $n$-dimensional systems, $\hat{K}$ mean-square stabilizes the systems if and only if $\rho(\Phi)<1$ where $\Phi\triangleq[A+(1-q)B\hat{K}]\otimes[A+(1-q)B\hat{K}]+(q-q^{2})(B\hat{K})\otimes(B\hat{K})$. This necessary and sufficient condition is used to search for the unstabilizable region (red region) in Fig. 4(a),4(b) with a step size $0.001$. The stabilizable regions (blue regions) in Fig. 4(a),4(b) are determined by Theorem 3.14 and 3.7, respectively. The error bounds provided by Theorem 3.7,3.14 for $B$-invertible systems under different $q$ are also presented in Fig. 5. The gray regions cannot be classified by Theorem 3.14 and 3.7. The conservatism curves are also shown in Fig. 6. Clearly, Theorem 3.14 exhibits less conservatism than Theorem 3.7 for this specific system, which makes the sample complexity based on (25) lower than that based on (23), as shown in Fig. 4(c),4(d).
Figure 5: Bounds on the estimation error $q-\hat{q}$ provided in Theorem 3.7,3.11,3.14 for Example 6.30,6.31.
Figure 6: Conservatism of Theorem 3.7,3.11,3.14 refers to the extent to which the minimum stabilizable $\hat{q}$ given by these theorems exceeds that given by the necessary and sufficient conditions.
7 Conclusion
This paper investigates the LQR problem over an unknown Bernoulli packet loss channel, where the unknown probability is estimated using finite samples. We analyze the maximum estimation error and the minimum sample size for the estimated controller to stabilize the system, and the performance gap between the estimated controller and the optimal controller. In the future, this work can be extended to the contexts of LQG and Markov packet loss channels. Furthermore, we aim to design an online controller, where the estimated probability can be improved over time, and analysis its performance.
[…]